OpenStudy (anonymous):

Find an equation of the tangent line to y = f(x) at x = a. Graph y = f(x) and the tangent line, to verify that you have the correct equation. f(x) = x^2 - 2, a = 1.

OpenStudy (anonymous):

first find f'(x) at x=a

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

it would be f'(a)=2a line equation: f(x)-f(x_0)=b(x-x_0) in this casen x_0=a, and b= 2a So: f(x)-a^2+2=2a(x-a), or writing it other way: y-a^2+2=2a(x-a) reoganizing a bit: y=2ax+(a^2-2-2a^2) , y=2ax-(a^2+2)

OpenStudy (anonymous):

if a=1, y=2x-3

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so how would i graph it?

OpenStudy (anonymous):



OpenStudy (anonymous):

ok the dot is at -3?

OpenStudy (anonymous):